The longer leg of a right triangle is
3cm
longer than the shorter leg. The hypotenuse is
6cm
longer than the shorter leg. Find the side lengths of the triangle.

Answer by mananth(16946)

short leg be x
The longer leg of a right triangle is
3cm longer than the shorter leg. (x+3)

The hypotenuse is 6cm longer than the shorter leg. (x+6)

x^2+(x+3)^2= (x+6)^2
x^2+x^2+6x+9 = x^2+12x+36
x^2-6x-27=0
x^2-9x+3x-27=0
x(x-9)+3(x-9)=0
(x-9)(x+3)=0
x=9
9,12, 15 cm are the lengths